# DecoderTracker: Decoder-Only End-To-End method for Multiple-Object Tracking

### v5

@@ 3.2.1 Fixed-Size Query Memory
FSQM maintains fixed-dimensional queries $q_{tr}^{\mathcal{M}}\in\mathbb{R}^{N\times d}$, where $N$ denotes the preset maximum number of tracking targets and $d$ represents the feature dimension. Compared to dynamic query mechanisms, this approach introduces:
$\hat{\omega}_{tr}^{\mathcal{M}}=(\underbrace{q_{tr}^{\mathcal{M}}}_{\text{static queries}},\underbrace{C_{\text{tr}}\in\mathbb{R}^{N}}_{\text{confidencevector}},\underbrace{\mathcal{ID}}_{\text{global ID pool}},\underbrace{BBox^{\mathcal{M}}}_{\text{bounding box}})$ (5)
to model memory for the continuous tracking process. During initialization, set $q_{ftr}^{\mathcal{M}}=\mathbf{0}$ and $C_{\text{tr}}^{(0)}=\mathbf{0}$ for all queries, where all queries are initialized to zero to avoid interference with networks that do not rely on attention mechanisms. Additionally, all inactive queries are also initialized to zero. Moreover, a global ID pool, denoted as $\mathcal{ID}$, is introduced. Each active tracking query is assigned a unique ID that accompanies the query as it moves across video frames. When a query is no longer associated with a tracked object, its corresponding ID is deleted and will not be reused. A new ID is assigned to new active queries. By the way, all inactive queries are assigned an ID of -1 and their bounding boxes are initialized with zero values.
During each frame processing, iterative updates are performed as follows:
@@
    where $S(\cdot)$ represents the prediction score function. The decay ensures that the relevance of older queries decreases over time, while new, active queries retain higher confidence scores.
It is also important to emphasize that, aside from this, the processing of queries in FSQM is consistent with that in DecoderTracker. The specific process will not be repeated here, but can be referred to in the previous sections.
@@
7:        $C_{\text{tr}}[k]\leftarrow c_{i}$Reset confidence to 1.0}
@@
3.2.2 Attention Mechanism Optimization
@@ 4.1 Datasets and Metrics
To validate DecoderTracker and DecoderTracker+’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[37], MOT17[38], and KITTI[39]. DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures. Comprising 100 videos featuring diverse dance styles, it underscores the significance of motion analysis in multi-object tracking.
MOT17 focuses on multiple object tracking in public spaces, primarily pedestrians. The dataset includes seven scenes, divided into training and testing clips, providing object detection annotations suitable for both online and offline tracking approaches.
KITTI serves as a research dataset for autonomous driving, encompassing stereo, optical flow, odometry, 3D object detection, and tracking tasks. Captured across diverse environments using advanced sensors, it provides raw data and task-specific benchmarks.
Evaluation Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA[40, 41]. These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance. HOTA assessed the tracking accuracy based on the spatial and temporal overlap between predicted and ground truth tracks. AssA measured the quality of associations between objects, while IDF1 quantified the accuracy of identity tracking. MOTA offered an overall evaluation of tracking accuracy, considering false positives, false negatives, and identity switches. These metrics formed the basis for a thorough evaluation of our method’s tracking capabilities.
@@ 4.2 Implementation Details
The majority of experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU. For equitable comparison with MOTR, training on the MOT17 dataset was performed on a single NVIDIA GeForce RTX 2080ti. Additionally, to better contrast with the MOTR series in terms of speed, inference speed tests were carried out on a single NVIDIA V100 GPU. One more thing, the image size for the input network is $640\times 640$.
@@
Figure 7: IDF1-MOTA-HOTA and IDF1-MOTA-FPS comparisons of state-of-the-art trackers with our proposed DecoderTracker and DecoderTracker+ on Dancetrack test sets. For the left image, the horizontal axis is MOTA, the vertical axis is IDF1, and the radius of the circle is HOTA. For the right image, the horizontal axis is FPS, the vertical axis is IDF1, and the radius of the circle is HOTA. ”*” represents additional data was used for training. Our DecoderTracker+* achieves the best IDF1 and HOTA and comparable speed performance.
Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[37] test set. Numbers are highlighted in bold when they correspond to either DecoderTracker or the superior metric from MOTR. * means that extra data was used during training.
Methods IDF1 HOTA AssA DetA MOTA
Non-End-to-end:
@@
Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17 [38] set. * means that extra data was used during training.
method IDF1 HOTA DetA AssA MOTA
Non-End-to-end:
YOLOTracker[43] 65.1 - - - 71.9
FairMOT[44] 72.3 59.3 60.9 58.0 73.7
MSPNet[45] 74.6 - - - 71.9
@@
DecoderTracker* 77.2 61.8 62.1 60.1 76.8
DecoderTracker+* 77.5 62.2 62.5 60.7 76.7
During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1.
For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [47], employing random displacement to generate video segments with pseudo-trajectories. Training comprised 120 epochs in the first stage, followed by 30 epochs in the second. The third stage extended to 20 epochs with TBSP, or 55 without. And, to address the issue of the limited scale of the MOT17 dataset, this paper employs a hybrid dataset. It integrates the MOT20 dataset and the complete CrowdHuman dataset, which has been spatially shifted, to expand data diversity. The amount of new data significantly exceeds the benchmark set (by approximately 4.3 times). To balance training bias, a stratified sampling mechanism was designed: within each epoch, samples from the MOT17 dataset are resampled twice, while samples from the supplementary sets are sampled once. This approach not only alleviates data skewness but also enhances the learning of benchmark set features, ensuring that the model maintains its original performance advantages while improving its adaptability to dense scenes.
4.3 Main Result
The comprehensive benchmarking in Table 1 demonstrates DecoderTracker’s dual superiority in training efficiency and inference speed over MOTR-series methods. This design achieves 68.7 training hours on a single GPU, equivalent to 7.1× higher GPU-hour efficiency versus MOTR’s 768 GPU-hour consumption (8 GPUs × 96 hours). The phased training regimen—12.1 hours for feature embedding, 30.8 hours for temporal modeling, and 43.4 hours for TBSP-enhanced prediction—systematically constructs tracking capabilities while maintaining 83% shorter wall-clock time than competitors. Inference performance reveals more pronounced advantages: DecoderTracker+ attains 28.8 FPS, 4.1× faster than MOTRv2’s 6.9 FPS,
For DanceTrack ,the experimental results shown in Table 2 reveal that DecoderTracker and its enhanced variant DecoderTracker+ achieve comparable precision levels across core evaluation metrics. Both models demonstrate superior detection accuracy over existing end-to-end methods, with DetA scores of 78.7 and 79.0 respectively, significantly outperforming MOTR’s 73.5. Their MOTA metrics (87.8 and 88.1) approach SOTA non-end-to-end methods like ByteTrack (89.5), while maintaining balanced HOTA performance at 54.2 and 54.5. This parity suggests that architectural refinements between the two variants primarily yield marginal improvements rather than fundamental performance shifts. Notably, both configurations exhibit identical IDF1 scores (54.2 and 54.4), confirming consistent identity preservation capabilities. The minimal performance delta between variants implies optimized parameter utilization rather than architectural overhauls. When extended with supplementary training data (DecoderTracker* variants), both models achieve substantial gains, attaining 91.1-91.9 MOTA and 60.8-61.2 HOTA. While their AssA scores (35.6-35.7) remain below MOTR’s 40.2, the detection-centric advantages manifest through DetA superiority over all end-to-end and most non-end-to-end competitors. These findings position the DecoderTracker framework as a robust baseline for end-to-end tracking, achieving detection reliability comparable to sophisticated multi-stage systems while preserving model efficiency.
In the MOT17 benchmark (Table 3), DecoderTracker shows strong performance, especially in complex tracking scenarios. While slightly behind MOTR in HOTA and AssA, it compensates with comparable DetA and excels in IDF1 by nearly 10 points. This advantage is due to its simplified network structure, which achieves competitive accuracy with less training data. Moreover, DecoderTracker maintains robust overall performance, as shown by its high MOTA score, making it suitable for real-world applications. When additional training data is introduced, DecoderTracker’s performance further improves. For instance, DecoderTracker* achieves a significant boost in IDF1 to 77.2, HOTA to 61.8, and MOTA to 76.8. Similarly, DecoderTracker+* also shows substantial gains. This clearly demonstrates the positive impact of more training data on DecoderTracker’s tracking capabilities, highlighting its potential for even better performance with increased data support.
@@ 4.3 Main Result
Table 4: Performance comparison between DecoderTracker and existing methods on the KITTI [39] set.
@@
Tracker HOTA MOTA AssA HOTA MOTA AssA
Non-End-to-end:
@@
Table 4 provides an overview of DecoderTracker’s performance on the KITTI benchmark. DecoderTracker achieves competitive results, with an HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively. These results position DecoderTracker as a strong contender in MOT on the KITTI.
@@ 4.4.1 More Discusstion About TBSP
The initial reason for developing the DecoderTrack method was not due to slow inference speed, but rather the extensive training resources required. Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs. Each epoch took around 0.42 hours, with each GPU processing about 3,000 images per hour.
@@
We believe that the TBSP, a weak supervision process, enables the method to preliminarily learn object motion information that would otherwise require many more training iterations. It’s akin to learning to ride a bicycle where training wheels help beginners to ride without obstacles while also giving them a feel for biking, which aids in eventually mastering the skill. A similar situation may exist here.
@@ 4.5.1 Experimental Analysis of FENet
Furthermore, we have made some modifications to the C2F module in YOLOv8, renaming our modified module to C2M. This change is specifically reflected in the BottleNeckF module. In YOLOv8, there is a BottleNeck module that differs slightly between the neck and backbone; we eliminated this difference and replaced the 3x3 convolution in the middle with a 1x1 convolution to further reduce the computational load. These modifications are illustrated in Figure 9, and their impact on performance is shown in Table 8.
@@ 5 Conclusion and Limitation
This paper proposes an efficient decoder-only end-to-end MOT method, which achieves faster inference speed and shorter training time compared to the baseline MOTR. Additionally, we introduce new training strategies aimed at accelerating the convergence of the method. Due to hardware limitations, we were unable to verify our training strategies on MOTR, but we believe these strategies are generalizable. Finally, we conducted a detailed and in-depth analysis of the slow inference process in MOTR, finding that dynamic data significantly impacts inference speed. Subsequently, we applied FSQM to address this issue, greatly enhancing the inference speed of the tracking-by-query end-to-end paradigm. Overall, this study provides a more efficient benchmark method for the multi-object tracking field, lays the foundation for the emergence of real-time end-to-end MOT methods, and offers insights for future development. In future work, we should further explore more efficient network architectures to build real-time MOT methods. Of course, we also plan to integrate the technologies from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to improve tracking performance.
Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection methods. Both methods fail under similar circumstances, as illustrated in Figure.10. This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the method. Clearly, despite this limitation, DecoderTracker still represents a significant improvement over MOTR. Although we attempted to use a mixed dataset to address this issue, our method still underperforms compared to SOTA tracking-by-detection methods. How to resolve this issue will also be a key focus of future research.

### v6

@@ 3.2.1 Fixed-Size Query Memory
FSQM maintains fixed-dimensional queries $q_{tr}^{\mathcal{M}}\in\mathbb{R}^{N_{m}\times d}$, where $N_{m}$ denotes the preset maximum number of tracking targets and $d$ represents the feature dimension. Compared to dynamic query mechanisms, this approach introduces:
$\hat{\omega}_{tr}^{\mathcal{M}}=(\underbrace{q_{tr}^{\mathcal{M}}}_{\text{static queries}},\underbrace{C_{\text{tr}}\in\mathbb{R}^{N_{m}}}_{\text{confidence vector}},\underbrace{\mathcal{ID}}_{\text{global ID pool}},\underbrace{BBox^{\mathcal{M}}}_{\text{bounding box}})$ (9)
to model memory for the continuous tracking process. During initialization, set $q_{ftr}^{\mathcal{M}}=\mathbf{0}$ and $C_{\text{tr}}^{(0)}=\mathbf{0}$ for all queries, where all queries are initialized to zero to avoid interference with networks that do not rely on attention mechanisms. Additionally, all inactive queries are also initialized to zero. Moreover, a global ID pool, denoted as $\mathcal{ID}$, is introduced. Each active tracking query is assigned a unique ID that accompanies the query as it moves across video frames. When a query is no longer associated with a tracked object, its corresponding ID is deleted and will not be reused. A new ID is assigned to new active queries. By the way, all inactive queries are assigned an ID of -1 and their bounding boxes are initialized with zero values. And the maximum number of tracking targets $N_{m}$ is set to be equal to the number of detect queries. Theoretically, as long as the actual number of tracked objects does not exceed this preset $N_{m}$, tracking performance remains unaffected. Based on empirical considerations, we thus set $N_{m}$ to match the number of object queries. During each frame processing, iterative updates are performed as follows:
@@
    The decay ensures that the relevance of older queries decreases over time, while new, active queries retain higher confidence scores.
It is also important to emphasize that, beyond the fixed-size query management (via injection/removal mechanisms) and the unique attention masking strategy of FSQM, the core handling of queries remains consistent with DecoderTracker: including the interaction between detect queries ($q_{d}$) and track queries ($q_{tr}$) in the decoder, cross-attention with FENet-derived feature maps, track query updates via TAN, and prediction of bounding boxes and object classes based on query hidden states.
@@
7:        $C_{\text{tr}}[k]\leftarrow 1$Reset confidence to 1.0}
@@
3.2.2 Attention Masking Strategy
@@ 4.1 Datasets and Metrics
Datasets: To validate the performance of DecoderTracker and DecoderTracker +, we evaluated the method on three demanding datasets: DanceTrack[43], MOT17[44], and KITTI[45]. DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures. Comprising 100 videos featuring diverse dance styles, it underscores the significance of motion analysis in multi-object tracking. MOT17 focuses on the tracking of multiple objects in public spaces, primarily by pedestrians. The dataset includes seven scenes, divided into training and testing clips, providing object detection annotations suitable for both online and offline tracking approaches. KITTI covers stereo, flow, odometry, detection and tracking; we use its multi-object tracking split for evaluation.
Evaluation Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation include HOTA, AssA, DetA ,IDF1 and MOTA[46, 47]. These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance. HOTA assessed the tracking accuracy based on the spatial and temporal overlap between predicted and ground truth tracks. AssA measured the quality of associations between objects, while IDF1 quantified the accuracy of identity tracking. MOTA offered an overall evaluation of tracking accuracy, considering false positives, false negatives, and identity switches. These metrics formed the basis for a thorough evaluation of our method’s tracking capabilities.
@@ 4.2 Implementation Details
The majority of experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU. For equitable comparison with MOTR, training on the MOT17 dataset was performed on a single NVIDIA GeForce RTX 2080ti. Additionally, to better contrast with the MOTR series in terms of speed, inference speed tests were carried out on a single NVIDIA V100 GPU. Furthermore, the image size for the input network is $640\times 640$. The number of object queries is configured as 60 for DanceTrack and KITTI and 300 for MOT17. These values are in line with the previous works [35, 22]. Consistently, the preset maximum number of tracking targets in the FSQM system is also aligned with these query numbers.
@@
Figure 7: IDF1-MOTA-HOTA and IDF1-FPS-HOTA comparisons of state-of-the-art trackers with our proposed DecoderTracker and DecoderTracker+ on DanceTrack test sets. For the left image, the horizontal axis is MOTA, the vertical axis is IDF1, and the radius of the circle is HOTA. For the right image, the horizontal axis is FPS, the vertical axis is IDF1, and the radius of the circle is HOTA. ”*” represents additional data was used for training. Our DecoderTracker+* achieves the best IDF1 and HOTA and comparable speed performance.
During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1. The random seed was fixed at 42 throughout all experiments to ensure reproducibility.
For MOT17, following MOTR’s data augmentation strategy [2], we incorporated additional data from CrowdHuman [48]. To convert static images into video-like sequences, we applied random affine transformations (primarily random translation and rotation) to individual images across simulated frames. Training comprised 120 epochs in the first stage, followed by 30 epochs in the second. The third stage extended to 20 epochs with TBSP, or 55 without. And, to address the issue of the limited scale of the MOT17 dataset, this paper employs a hybrid dataset. It integrates the MOT20 dataset and the complete CrowdHuman dataset, which has been spatially shifted, to expand data diversity. The amount of new data significantly exceeds the benchmark set (by approximately 4.3 times). To balance training bias, a stratified sampling mechanism was designed: within each epoch, samples from the MOT17 dataset are resampled twice, while samples from the supplementary sets are sampled once. This approach not only alleviates data skewness but also enhances the learning of benchmark set features, ensuring that the model maintains its original performance advantages while improving its adaptability to dense scenes.
4.3 Main Result
The comprehensive benchmarking in Table 1 demonstrates DecoderTracker’s dual superiority in training efficiency and inference speed over MOTR-series methods. This design achieves 68.7 training hours on a single GPU, equivalent to 7.1× higher GPU-hour efficiency versus MOTR’s 768 GPU-hour consumption (8 GPUs × 96 hours). The phased training regimen—12.1 hours for feature embedding, 30.8 hours for temporal modeling, and 43.4 hours for TBSP-enhanced prediction—systematically constructs tracking capabilities while maintaining 83% shorter wall-clock time than competitors. Inference performance reveals more pronounced advantages: DecoderTracker+ attains 28.8 FPS, 4.1× faster than MOTRv2’s 6.9 FPS.
Table 2: Performance comparison between DecoderTracker and existing methods on the DanceTrack [43] test set. Bold numbers indicate either DecoderTracker results or superior MOTR metrics, while our method uses color highlighting for distinction. Asterisk (*) denotes methods using extra training data.
Methods Metrics
IDF1 HOTA AssA DetA MOTA
Non-End-to-end
@@ 4.3 Main Result
As evidenced in Table 2, DecoderTracker establishes new performance benchmarks for end-to-end multi-object tracking on the challenging DanceTrack benchmark. Our method demonstrates superior detection capabilities, with DecoderTracker and DecoderTracker+ achieving substantially higher DetA scores than all existing end-to-end approaches. This detection advantage directly translates to competitive MOTA performance that approaches state-of-the-art non-end-to-end methods, underscoring the efficacy of our detection-centric architecture. Notably, both variants maintain balanced overall tracking performance comparable to the strongest end-to-end baseline while exhibiting consistent identity preservation metrics (IDF1). The minimal performance variance between DecoderTracker and its enhanced version suggests optimized architectural efficiency rather than fundamental design differences. When augmented with additional training data, our framework achieves transformative performance gains, matching or exceeding leading non-end-to-end methods across multiple metrics. These data-enhanced variants particularly showcase exceptional MOTA and IDF1 results, demonstrating our architecture’s scalability and untapped potential. While association accuracy (AssA) remains an area for future improvement relative to specialized baselines, the consistent DetA superiority highlights our method’s core strength in target localization. This positions DecoderTracker as a robust foundational framework that bridges the performance gap between end-to-end paradigms and complex multistage trackers while maintaining model efficiency.
Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17 [44] test set. Bold numbers indicate either DecoderTracker results or superior MOTR metrics, while our method uses color highlighting for distinction. Asterisk (*) denotes methods using extra training data.
Methods Metrics
IDF1 HOTA DetA AssA MOTA
Non-End-to-end
MSPNet [30] 74.6 71.9
@@
DQFormer [26] 71.6 59.2 60.6 58.1 75.2
DecoderTracker* 73.3 61.7 62.5 60.4 76.7
DecoderTracker+* 73.4 62.0 62.1 60.2 77.2
In the MOT17 benchmark (Table 3), DecoderTracker exhibits strong performance, with a distinct advantage in IDF1 despite marginal deficits relative to MOTR in HOTA and AssA. Its simplified architecture achieves competitive accuracy with reduced training data while maintaining robust overall performance (high MOTA) suitable for real-world applications. Augmented training data further enhances its capabilities: DecoderTracker* and DecoderTracker+* demonstrate significant improvements in IDF1, HOTA, and MOTA, underscoring the positive impact of increased data volume. The slight gaps in HOTA and AssA compared to MOTR can be derived from two factors: first, the streamlined architecture (encoder removed) affects the modeling of long-term temporal correlations, which is critical for HOTA and AssA, which are sensitive to inter-object association accuracy; second, the weak supervision based on TBSP in the three-stage training strategy constrains the precision of the association in complex scenarios, in contrast to the more aggressive TALA strategy of MOTR. To address this, we propose three potential future improvements: integrating the release-fetch supervision from MOTRv3 and the coopetition label assignment strategy from CO-MOT to balance the label assignment weights between detection and tracking queries; enhancing the temporal modeling capability of the TAN module by introducing a dynamic adjustment mechanism for cross-frame attention; and optimizing the matching strategy between static queries and dynamic objects in FSQM to reduce interference from invalid queries on association accuracy. It should be noted that since the focus of this paper is to verify the efficiency advantages of the decoder-only architecture and the effectiveness of the basic training strategy, the improvement directions above will be explored in subsequent studies. Due to the constraints in length and scope of the current research, in-depth exploration of these directions is not included in this paper.
@@
Table 4: Performance comparison between DecoderTracker and existing methods on the KITTI [45] test set. Bold numbers indicate the best performance in each category.
@@
Methods HOTA MOTA AssA HOTA MOTA AssA
Non-End-to-end
@@
Table 4 provides an overview of DecoderTracker’s performance on the KITTI benchmark. DecoderTracker achieves competitive results, with HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively. These results position DecoderTracker as a strong contender in MOT on the KITTI.
@@ 4.4.1 More Discusstion About TBSP
The initial reason for developing the DecoderTracker method was not due to slow inference speed, but rather the extensive training resources required. Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs. Each epoch took around 0.42 hours, with each GPU processing about 3,000 images per hour.
@@ 4.5.1 Experimental Analysis of FENet
Furthermore, we have made some modifications to the C2F module in YOLOv8, renaming our modified module to C2M. This change is specifically reflected in the BottleNeckF module. In YOLOv8, there is a BottleNeck module that differs slightly between the neck and backbone; we eliminated this difference and replaced the 3x3 convolution in the middle with a 1x1 convolution to further reduce the computational load. These modifications, grounded in YOLOv8’s proven stability and migration performance (with YOLO11’s key changes focusing more on the head section, less relevant to our feature extraction adjustments), are illustrated in Figure 9, and their impact on performance is shown in Table 8.
@@ 5 Conclusion and Limitation
This paper presents DecoderTracker and its enhanced variant DecoderTracker+, two novel end-to-end MOT methods based on a decoder-only architecture, aiming to address the issues of high computational resource consumption, suboptimal inference speed, and prolonged training time in traditional transformer-based end-to-end MOT methods like MOTR. DecoderTracker optimizes the network architecture by removing the redundant encoder and adopting a lightweight Feature Extraction Network (FENet) derived from YOLOv8, which enhances efficiency and inference speed. It also introduces a three-stage training strategy, incorporating TBSP as a weakly supervised training step to accelerate convergence and reduce training time. To tackle optimization challenges from dynamic data, DecoderTracker+ is proposed by integrating FSQM and refining attention layers with an attention masking strategy. FSQM maintains a constant number of tracking queries, enabling better compatibility with modern deep learning frameworks and facilitating deployment, while the masking strategy ensures that inactive queries do not interfere with computations. Experimental results on benchmarks including DanceTrack, MOT17, and KITTI demonstrate that both methods outperform MOTR without additional tricks. DecoderTracker+ achieves an inference speed of 28.8 FPS, which is 2-3 times faster than MOTR, while maintaining competitive performance in metrics such as HOTA, IDF1, and MOTA. This work provides an efficient baseline for end-to-end MOT, laying the foundation for real-time end-to-end MOT methods and offering insights for engineering deployment, with potential future improvements including enhancing temporal modeling and optimizing label assignment strategies .
DecoderTracker’s primary limitation remains its accuracy gap on MOT17 compared to TBD(e.g., Deep OC-SORT), similar to MOTR’s challenge. We attribute this in part to the limited volume of MOT17 training data, which affects end-to-end training effectiveness (Figure 10). Although our method significantly improves speed and outperforms MOTR, it still trades accuracy for real-time efficiency, a core design choice. Additional challenges include robustness in high-occlusion scenarios (Fig. 10) and domain generalization. Future work will focus on improving occlusion handling, exploring domain adaptation, and refining label assignment strategies to close the accuracy gap.